OpenStudy (anonymous):

Could anyone explain to me how to write a function rule for this table, as well as walk me through on how to write the actual function rule? x 2 4 6 y 1 0 –1

OpenStudy (anonymous):

I noticed that every time x increases by 2, y decreases by 1. So, my first guess was y = -1/2 x, but I am still not completely sure..

OpenStudy (anonymous):

well you have the slope correct, now just add the y-intercept to y = 1/2 x

OpenStudy (anonymous):

Would it be y = 2 - 1/2 x ?

OpenStudy (anonymous):

looks right to me!

OpenStudy (anonymous):

Great! Thank you very much for the assistance. :)